# RSA: Resolving Scale Ambiguities in Monocular Depth Estimators through Language Descriptions

### v1

@@ 4 Experiments
Models Scaling Dataset $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ Abs Rel $\downarrow$ $\text{RMSE}_{\log}\downarrow$ RMSE $\downarrow$ ZoeDepth Image KITTI 0.971 0.996 0.999 0.054 0.082 2.281 Monodepth2 Median KITTI 0.877 0.959 0.981 0.115 0.193 4.863 ZeroDepth DA - 0.892 0.961 0.977 0.102 0.196 4.378 ZeroDepth DA,Median - 0.886 0.965 0.984 0.105 0.178 4.194 DPT Median KITTI 0.950 0.994 0.999 0.069 0.100 3.365 Linear fit KITTI 0.974 0.997 0.999 0.052 0.080 2.198 Global KITTI 0.959 0.995 0.999 0.062 0.092 2.575 Image KITTI 0.961 0.995 0.999 0.064 0.092 2.379 Image NYUv2,KITTI 0.956 0.989 0.993 0.066 0.098 2.477 RSA (Ours) KITTI 0.963 0.995 0.999 0.061 0.090 2.354 RSA (Ours) NYUv2,KITTI 0.962 0.994 0.998 0.060 0.089 2.342 MiDas Median KITTI 0.856 0.959 0.988 0.138 0.204 6.372 Linear fit KITTI 0.824 0.952 0.989 0.154 0.174 3.833 Global KITTI 0.729 0.939 0.978 0.192 0.212 4.811 Image KITTI 0.749 0.949 0.982 0.164 0.199 4.254 Image NYUv2,KITTI 0.718 0.943 0.979 0.171 0.211 4.456 RSA (Ours) KITTI 0.798 0.948 0.981 0.163 0.185 4.082 RSA (Ours) NYUv2,KITTI 0.782 0.946 0.980 0.160 0.194 4.232 DepthAnything Median KITTI 0.925 0.986 0.996 0.091 0.129 3.648 Linear fit KITTI 0.824 0.896 0.922 0.149 0.224 3.595 Global KITTI 0.663 0.932 0.981 0.191 0.228 5.273 Image KITTI 0.768 0.951 0.983 0.162 0.195 4.483 Image NYUv2,KITTI 0.697 0.933 0.977 0.181 0.218 4.824 RSA (Ours) KITTI 0.780 0.958 0.988 0.160 0.189 4.437 RSA (Ours) NYUv2,KITTI 0.756 0.956 0.987 0.158 0.191 4.457
Table 2: Quantitative results on KITTI Eigen Split. Global refers to a single scale and shift for entire dataset. Image denotes predicting scale and shift from an image. Red denotes scaling that uses ground truth. Median scaling per image is a common practice for evaluation. Linear fit denotes fitting to ground truth for each image. DA refers to domain adaptation. ZoeDepth performs per-pixel refinement. RSA (yellow) when trained with NYUv2 and KITTI generalizes better than using image to predict the transformation parameters.
@@
Zero-shot Generalization. Considering the smaller domain gap in language descriptions across various scenes, we perform a zero-shot transfer experiment to demonstrate RSA’s generalization ability. We evaluate the models on the Sun-RGBD without fine-tuning. As shown in Table 3, RSA achieves superior results compared to baselines, existing methods, and ground truth scaling. This suggests that language descriptions offer stable scaling for generalization across diverse data domains. Note that a single global scale and shift are ineffective for both indoor and outdoor settings. Therefore, for the “Global" model, we fit it only to NYUv2 to obtain a reasonable global scale and shift for the zero-shot Sun-RGBD evaluation.
@@
Models Scaling Dataset $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ Abs Rel $\downarrow$ $\log_{10}\downarrow$ RMSE $\downarrow$ Adabins - NYUv2 0.771 0.944 0.983 0.159 0.068 0.476 DepthFormer - NYUv2 0.815 0.970 0.993 0.137 0.059 0.408 ZoeDepth-X Image NYUv2 0.857 - - 0.124 - 0.363 ZoeDepth-M12 Image NYUv2 0.864 - - 0.119 - 0.346 ZoeDepth-M12 Image NYUv2, KITTI 0.856 - - 0.123 - 0.356 DPT Linear Fit SUN-RGBD 0.812 0.967 0.993 0.139 0.059 0.412 Global NYUv2 0.773 0.945 0.984 0.154 0.071 0.482 Image NYUv2,KITTI 0.778 0.953 0.984 0.153 0.068 0.478 RSA (Ours) NYUv2,KITTI 0.781 0.953 0.986 0.152 0.066 0.463 MiDas Linear Fit SUN-RGBD 0.632 0.912 0.971 0.241 0.102 1.132 Global NYUv2 0.572 0.889 0.956 0.297 0.132 1.464 Image NYUv2,KITTI 0.594 0.895 0.962 0.275 0.125 1.374 RSA (Ours) NYUv2,KITTI 0.612 0.903 0.964 0.268 0.122 1.302 DepthAnything Linear Fit SUN-RGBD 0.878 0.979 0.995 0.113 0.054 0.332 Global NYUv2 0.534 0.872 0.951 0.313 0.138 1.692 Image NYUv2,KITTI 0.588 0.892 0.963 0.279 0.126 1.392 RSA (Ours) NYUv2,KITTI 0.621 0.915 0.970 0.238 0.099 1.024
@@
Table 3: Zero-shot generalization to SUN-RGBD. The models are tested on the Sun-RGBD without any fine-tuning. For ZoeDepth, X indicates no pre-training, and M12 indicates 12 datasets for pre-training. ZoeDepth results were taken from their original manuscripts, using a depth decoder for scaling.
@@
Prompt design for input text. In Table 4 we investigate different designs of RSA text prompts in training and how they affect the performance. Here to make the experiment more controllable, we use only structural text and only produce one caption for each image to train each model.
From the 1st and 2nd rows, we use an object detector Detic [86] to produce only foreground objects in images and form input text using only foreground objects. We observe that if the input text only specifies the types of objects and their numbers, RSA can still accurately predict the scale for indoor scenes, as these spaces are typically filled with various pieces of furniture. However, this approach performs poorly for outdoor scenes, which tend to be more open and sparse. For instance, parking lots of different sizes may contain varying numbers of cars: a small lot may be crowded, while a large lot may appear empty. However, from the 3rd and 4th rows, where we use text formed using segmentation results, we observe that after including backgrounds, the model works better especially for outdoor scenes, since different backgrounds (like wall, sidewalk, highway, road, sky, house, building) can reflect different scales.
@@
Figure 5: Left: Predicted inverse scale w.r.t. median depth ground truth. For combined KITTI and NYUv2 with DepthAnything model, we fit an inverse proportional function for the predicted inverse scale in the test set (each point is an image), to verify that the scale is proportional to the median depth. It shows that larger scenes can be predicted with larger scales. Two clusters could be noticed since cross-datasets have domain gaps in their depth distribution. The fitted function is $y=\frac{1}{7.65x+12.06}$.
@@ 5 Discussion
Conclusion. We present the first study exploring whether language as an additional input modality can resolve the scale ambiguity in monocular depth estimation, an issue particularly relevant in the context of the recent trend towards large-scale mixed dataset training. We propose a framework, RSA, which learns to convert a scaleless (relative) depth map to metric depth using language input. RSA utilizes the pre-trained CLIP encoder, and maps a language description of the image to scale and shift factors that transform the relative depth to metric depth. RSA is validated through extensive experiments on two benchmark datasets and three pre-trained relative depth models, demonstrating significant promise by drastically closing the gap between depth accuracy and its theoretical upper bound (that relies on ground truth), validating the hypothesis that language carries valuable scale information that could be used to enhance depth estimation. Moreover, we demonstrate that a unified model trained on both indoor and outdoor datasets with diverse scales generalizes across both scenarios, highlighting the robustness of language information in inferring scales. Finally, a zero-shot transfer experiment shows that the minimal domain gap of language description across scenes further generalizes to unseen data domains, without additional training. The generalizability of RSA stems of our choice of modality, language, which is invariant to nuisance variability that are present in images from lighting conditions and occlusions to specular reflectance and object deformations. Our results demonstrate that RSA is a viable choice to support relative to metric scale alignment for general-purpose monocular depth estimators.
Limitations and future work. Currently, although a simple global scale has proven effective, it may not always be sufficiently expressive for converting relative depth to absolute depth, especially when the relative depth is inaccurate. In such cases, global scaling may not adequately refine the depth map. Future research may include extending RSA to handle finer adjustments and investigate the potential of inferring region-wise or even pixel-wise scales using language input. Lastly, while language boasts high ease of use, RSA is also vulnerable to malicious users who may choose captions to steer predictions incorrectly.

### v2

@@ 4 Experiments
Models Scaling Dataset $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ Abs Rel $\downarrow$ $\text{RMSE}_{\log}\downarrow$ RMSE $\downarrow$ ZoeDepth Image KITTI 0.971 0.996 0.999 0.054 0.082 2.281 Monodepth2 Median KITTI 0.877 0.959 0.981 0.115 0.193 4.863 ZeroDepth DA - 0.892 0.961 0.977 0.102 0.196 4.378 ZeroDepth DA,Median - 0.886 0.965 0.984 0.105 0.178 4.194 DPT Median KITTI 0.950 0.994 0.999 0.069 0.100 3.365 Linear fit KITTI 0.974 0.997 0.999 0.052 0.080 2.198 Global KITTI 0.959 0.995 0.999 0.062 0.092 2.575 Image KITTI 0.961 0.995 0.999 0.064 0.092 2.379 Image NYUv2,KITTI 0.956 0.989 0.993 0.066 0.098 2.477 Image NYUv2,KITTI,VOID 0.952 0.987 0.993 0.068 0.098 2.568 RSA (Ours) KITTI 0.963 0.995 0.999 0.061 0.090 2.354 RSA (Ours) NYUv2,KITTI 0.962 0.994 0.998 0.060 0.089 2.342 RSA (Ours) NYUv2,KITTI,VOID 0.961 0.994 0.999 0.064 0.091 2.335 MiDas Median KITTI 0.856 0.959 0.988 0.138 0.204 6.372 Linear fit KITTI 0.824 0.952 0.989 0.154 0.174 3.833 Global KITTI 0.729 0.939 0.978 0.192 0.212 4.811 Image KITTI 0.749 0.949 0.982 0.164 0.199 4.254 Image NYUv2,KITTI 0.718 0.943 0.979 0.171 0.211 4.456 Image NYUv2,KITTI,VOID 0.683 0.931 0.972 0.165 0.232 4.862 RSA (Ours) KITTI 0.798 0.948 0.981 0.163 0.185 4.082 RSA (Ours) NYUv2,KITTI 0.782 0.946 0.980 0.160 0.194 4.232 RSA (Ours) NYUv2,KITTI,VOID 0.794 0.960 0.992 0.155 0.179 3.989 DepthAnything Median KITTI 0.925 0.986 0.996 0.091 0.129 3.648 Linear fit KITTI 0.824 0.896 0.922 0.149 0.224 3.595 Global KITTI 0.663 0.932 0.981 0.191 0.228 5.273 Image KITTI 0.768 0.951 0.983 0.162 0.195 4.483 Image NYUv2,KITTI 0.697 0.933 0.977 0.181 0.218 4.824 Image NYUv2,KITTI,VOID 0.678 0.924 0.974 0.186 0.243 5.021 RSA (Ours) KITTI 0.780 0.958 0.988 0.160 0.189 4.437 RSA (Ours) NYUv2,KITTI 0.756 0.956 0.987 0.158 0.191 4.457 RSA (Ours) NYUv2,KITTI,VOID 0.786 0.967 0.995 0.147 0.179 4.143
Table 2: Quantitative results on KITTI Eigen Split. RSA (yellow), especially when trained with multiple datasets, generalizes better than using images to predict the transformation parameters. Please refer to Table 1 for more details about notations.
Quantitative results. We show results on NYUv2 in Table 1, KITTI in Table 2, and VOID in Table 3, where we improve over baselines across all evaluation metrics and approach the performance of using ground truth for scaling. Following DPT, we optimize the scale and shift for the “Global” scaling baseline over the training set and use it for evaluation (i.e., same scale and shift for all test samples). We obtain the “Image” baseline by substituting CLIP text features with CLIP image features. Following  [57], we perform a linear regression to find the scale and shift that minimizes the least-square error between the ground truth metric depth and the predicted metric depth. Additionally, we also test median scaling [26], a common practice for evaluation, which uses the ratio between the median of depth prediction and ground truth as the scaling factor. Both linear fitting and median scaling are shown to demonstrate what is achievable if one were to directly fit to ground truth. We train separate RSA models for each dataset, as well as a unified RSA model combining both KITTI and NYUv2, or all KITTI, NYUv2, and VOID.
Although RSA trained on a single dataset may achieve slightly better performance than the unified model, the difference is minimal. In some metrics, the unified model even outperforms, demonstrating the generalizability of using language as input, given the narrow domain gap for language. Additionally, RSA models consistently outperform “Image” baselines in both in-domain and cross-domain scenarios and are comparable with existing methods under various settings. Our method significantly narrows the performance gap to that of using ground truth for scaling, validating the effectiveness of using language instead of the input image to predict scale.
To examine our scale predictions in more detail, Figure 3 shows a curve fitting plotted for our predicted (inverse) scale against the median of the ground truth. The trend line shows that the scale inferred from text descriptions matches well with median scaling, which is a robust estimator.
Models Scaling Dataset $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ Abs Rel $\downarrow$ $\log_{10}\downarrow$ RMSE $\downarrow$ DPT Median VOID 0.782 0.962 0.990 0.150 0.064 0.340 Global NYUv2 (zero-shot) 0.456 0.743 0.912 0.312 0.136 0.896 Image NYUv2,KITTI (zero-shot) 0.516 0.812 0.936 0.289 0.112 0.634 Image NYUv2,KITTI,VOID 0.534 0.827 0.941 0.266 0.108 0.545 RSA (Ours) NYUv2,KITTI (zero-shot) 0.601 0.886 0.970 0.254 0.096 0.444 RSA (Ours) NYUv2,KITTI,VOID 0.598 0.877 0.956 0.248 0.100 0.475 MiDas Median VOID 0.500 0.781 0.899 0.347 0.130 0.829 Global NYUv2 (zero-shot) 0.268 0.597 0.735 0.512 0.193 1.346 Image NYUv2,KITTI (zero-shot) 0.304 0.626 0.812 0.487 0.159 0.913 Image NYUv2,KITTI,VOID 0.389 0.743 0.911 0.392 0.139 0.652 RSA (Ours) NYUv2,KITTI (zero-shot) 0.392 0.696 0.892 0.448 0.148 0.660 RSA (Ours) NYUv2,KITTI,VOID 0.535 0.829 0.945 0.280 0.112 0.528 DepthAnything Median VOID 0.249 0.465 0.643 0.682 0.254 1.251 Global NYUv2 (zero-shot) 0.084 0.194 0.376 1.674 0.389 2.046 Image NYUv2,KITTI (zero-shot) 0.093 0.215 0.412 1.497 0.345 1.963 Image NYUv2,KITTI,VOID 0.323 0.612 0.768 0.589 0.196 0.956 RSA (Ours) NYUv2,KITTI (zero-shot) 0.104 0.262 0.450 1.287 0.323 1.716 RSA (Ours) NYUv2,KITTI,VOID 0.374 0.673 0.837 0.477 0.168 0.792
Table 3: Quantitative results on VOID. In zero-shot generalization and multi-dataset training (including the target dataset), RSA outperforms image scaling due to the robustness of text, which supports better generalization. Please refer to Table 1 for more details about notations.
[…]
Figure 3: Left: Predicted inverse scale w.r.t. median depth ground truth. Larger scenes tend to have larger median ground truth depth. For RSA trained on combined KITTI and NYUv2 with Depth Anything model, we fit an inverse proportional function for the predicted inverse scale in the test set (each point is an image), to verify that the scale is proportional to the median depth, that larger scenes are predicted with larger scales.
@@
Zero-shot Generalization. Considering the smaller domain gap in language descriptions across various scenes, we perform a zero-shot transfer experiment to demonstrate RSA’s generalization ability. We evaluate the models on the Sun-RGBD [62] and DDAD [28] without fine-tuning. As shown in Table 4 and Table 5, RSA achieves superior results compared to baselines, existing methods, and ground truth scaling. This suggests that language descriptions offer a viable option for relative to metric transfer when generalizing across diverse data domains. Note that a single global scale and shift are ineffective for both indoor and outdoor settings. Therefore, for the “Global” model, we fit it only to NYUv2 to obtain a reasonable global scale and shift for the zero-shot Sun-RGBD evaluation, and fit it to KITTI for DDAD evaluation.
@@
Models Scaling Dataset $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ Abs Rel $\downarrow$ $\log_{10}\downarrow$ RMSE $\downarrow$ Adabins - NYUv2 0.771 0.944 0.983 0.159 0.068 0.476 DepthFormer - NYUv2 0.815 0.970 0.993 0.137 0.059 0.408 ZoeDepth-X Image NYUv2 0.857 - - 0.124 - 0.363 ZoeDepth-M12 Image NYUv2 0.864 - - 0.119 - 0.346 ZoeDepth-M12 Image NYUv2, KITTI 0.856 - - 0.123 - 0.356 DPT Linear Fit SUN-RGBD 0.812 0.967 0.993 0.139 0.059 0.412 Global NYUv2 0.773 0.945 0.984 0.154 0.071 0.482 Image NYUv2,KITTI 0.778 0.953 0.984 0.153 0.068 0.478 RSA (Ours) NYUv2,KITTI 0.781 0.953 0.986 0.152 0.066 0.463 RSA (Ours) NYUv2,KITTI,VOID 0.788 0.953 0.986 0.150 0.065 0.458 MiDas Linear Fit SUN-RGBD 0.632 0.912 0.971 0.241 0.102 1.132 Global NYUv2 0.572 0.889 0.956 0.297 0.132 1.464 Image NYUv2,KITTI 0.594 0.895 0.962 0.275 0.125 1.374 RSA (Ours) NYUv2,KITTI 0.612 0.903 0.964 0.268 0.122 1.302 RSA (Ours) NYUv2,KITTI,VOID 0.623 0.908 0.968 0.253 0.116 1.223 DepthAnything Linear Fit SUN-RGBD 0.878 0.979 0.995 0.113 0.054 0.332 Global NYUv2 0.534 0.872 0.951 0.313 0.138 1.692 Image NYUv2,KITTI,VOID 0.588 0.892 0.963 0.279 0.126 1.392 RSA (Ours) NYUv2,KITTI 0.621 0.915 0.970 0.238 0.099 1.024 RSA (Ours) NYUv2,KITTI,VOID 0.645 0.927 0.978 0.203 0.095 1.137
@@
Table 4: Zero-shot generalization to SUN-RGBD. With more training datasets for scale prediction, RSA model generalizes better due to the robustness of text, but predicting scale using images suffers from domain gaps among training images. The models are tested on the Sun-RGBD without any fine-tuning. For ZoeDepth, X indicates no pre-training, and M12 indicates 12 datasets for pre-training. ZoeDepth results were taken from their original manuscripts, using a depth decoder for scaling. Please refer to Table 1 for detailed notations.
Models Scaling Dataset $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ Abs Rel $\downarrow$ $\text{RMSE}_{\log}\downarrow$ RMSE $\downarrow$ Adabins - KITTI 0.790 - - 0.154 - 8.560 NeWCRFs - KITTI 0.874 - - 0.119 - 6.183 ZoeDepth-X Image KITTI 0.790 - - 0.137 - 7.734 ZoeDepth-M12 Image KITTI 0.835 - - 0.129 - 7.108 ZoeDepth-M12 Image NYUv2, KITTI 0.824 - - 0.138 - 7.225 DPT Linear Fit DDAD 0.802 0.954 0.990 0.163 0.254 10.342 Global KITTI 0.752 0.925 0.969 0.183 0.312 15.967 Image NYUv2,KITTI 0.763 0.931 0.975 0.179 0.308 14.468 Image NYUv2,KITTI,VOID 0.731 0.910 0.962 0.191 0.324 16.132 RSA (Ours) NYUv2,KITTI 0.777 0.938 0.981 0.171 0.284 13.539 RSA (Ours) NYUv2,KITTI,VOID 0.768 0.942 0.983 0.165 0.276 12.437 MiDas Linear Fit DDAD 0.664 0.912 0.973 0.209 0.301 18.341 Global KITTI 0.603 0.864 0.925 0.253 0.336 20.594 Image NYUv2,KITTI 0.616 0.883 0.934 0.231 0.331 20.034 Image NYUv2,KITTI,VOID 0.564 0.862 0.925 0.243 0.352 22.689 RSA (Ours) NYUv2,KITTI 0.631 0.903 0.966 0.223 0.325 19.342 RSA (Ours) NYUv2,KITTI,VOID 0.642 0.908 0.966 0.218 0.331 18.293 DepthAnything Linear Fit DDAD 0.673 0.932 0.983 0.182 0.286 18.423 Global KITTI 0.612 0.883 0.963 0.221 0.323 21.345 Image NYUv2,KITTI 0.623 0.890 0.968 0.217 0.316 20.834 Image NYUv2,KITTI,VOID 0.586 0.874 0.956 0.243 0.348 22.351 RSA (Ours) NYUv2,KITTI 0.642 0.903 0.976 0.207 0.303 19.715 RSA (Ours) NYUv2,KITTI,VOID 0.648 0.905 0.975 0.198 0.297 18.984
Table 5: Zero-shot generalization to DDAD. With more training datasets for scale prediction, RSA model achieves a better generalization due to the robustness of text, but predicting scale using images suffers from domain gaps among training images. Models are tested on the DDAD without any fine-tuning. Please refer to Table 1 and Table 4 for more details about notations.
@@
Prompt design for input text. In Table 6, we investigate different designs of RSA text prompts in training and how they affect the performance. Here, to make the experiment more controllable, we use only structured text and only produce one caption for each image to train each model.
In the 1st and 2nd rows of Table 6, we use an object detector Detic [104] to produce only foreground objects in images and form input text using only foreground objects. We observe that if the input text only specifies the types of objects and their numbers, RSA can still accurately predict the scale for indoor scenes, as these spaces are typically filled with various pieces of furniture. However, this approach performs poorly for outdoor scenes, which tend to be more open and sparse. For instance, parking lots of different sizes may contain varying numbers of cars: a small lot may be crowded, while a large lot may appear empty. In the 3rd and 4th rows of Table 6, we use text formed using segmentation results; we observe that after including background classes, the model works better especially for outdoor scenes, since different backgrounds (like wall, sidewalk, highway, road, sky, house, building) can reflect different scales.
@@ 5 Discussion
Conclusion. We present the first study exploring whether language, as an additional input modality, can resolve the scale ambiguity in monocular depth estimation, an issue particularly relevant in the context of the recent trend towards large-scale mixed dataset training. We propose a framework, RSA, which learns to convert a scaleless (relative) depth map to metric depth using language descriptions as input. RSA utilizes the pre-trained CLIP encoder, and maps a language description of the image to scale and shift factors that transform the relative depth to metric depth. RSA is validated through extensive experiments on three benchmark datasets and three pre-trained relative depth models, demonstrating significant promise by drastically closing the gap between depth estimation accuracy and its upper bound (that relies on ground truth), validating the hypothesis that language carries valuable scale information that could be used to enhance depth estimation. Moreover, we demonstrate that a unified model trained on both indoor and outdoor datasets with diverse scene compositions generalizes across both scenarios, highlighting the robustness of language information in inferring scale. Finally, a zero-shot transfer experiment shows that the minimal domain gap of language description across scenes further generalizes to unseen data domains, without additional training. The generalizability of RSA stems of our choice of modality, language, which is invariant to nuisance variability that are present in images from lighting conditions and occlusions to specular reflectance and object deformations. Our results demonstrate that RSA is a viable choice to support relative to metric scale alignment for general-purpose monocular depth estimators.
Limitations and future work. We assume that the estimated 3D scene is up to an unknown scale. Although a simple global scale has proven effective, it may not always be sufficiently expressive for converting relative depth to absolute depth, especially when the relative depth is inaccurate. In such cases, global scaling may not adequately recover a high-fidelity metric-scaled depth map due to the presence of outliers. To address this, one may need to refine the relative depth outputted by general-purpose monocular depth estimators. Future research may include extending RSA to handle finer adjustments to also refine depth estimates and investigate the potential of inferring region-wise or even pixel-wise scales using language input. Lastly, while language boasts high ease of use, RSA is also vulnerable to malicious users who may choose captions to steer predictions incorrectly.
@@
Supplementary Material
A Evaluation Metrics
Following [9, 46, 96], we evaluate RSA and baseline methods quantitatively using mean absolute relative error (Abs Rel), root mean square error (RMSE), absolute error in log space $\left(\log_{10}\right)$, logarithmic root mean square error ($\text{RMSE}_{\log}$) and threshold accuracy $\left(\delta_{i}\right)$. The evaluation metrics are summarized in Table 8.
[…]
Table 8: Evaluation metrics. $y$ denotes predictions, $y^{*}$ denotes ground truth, $N$ denotes the valid number of pixels, $\Omega_{v}$ denotes the image space where the ground truth is valid, and $(i,j)$ denotes pixel coordinate,
B Datasets
We conduct experiments on five datasets in total. Details of each dataset are provided below.
KITTI [21, 23] contains 61 driving scenes with research in autonomous driving and computer vision. It contains calibrated RGB images with sychronized point clouds from Velodyne lidar, inertial, GPS information, etc. We used Eigen split [12], consisting of 23,488 training images and 697 testing images. We follow the evaluation protocol of [14] for our experiments.
VOID [73] comprises indoor (laboratories, classrooms) and outdoor (gardens) scenes with synchronized $640\times 480$ RGB images and with ground truth depth maps captured by active stereo. For the purpose of simulating sequences observed in visual inertial odometry (VIO), e.g., XIVO [16]), VOID contains 56 sequences with challenging 6 DoF motion captured on rolling shutter. 48 sequences (about 45,000 frames) are assigned for training and 8 for testing (800 frames). We follow the evaluation protocol of [73] where output depth is evaluated against ground truth within the depth range between 0.2 and 5.0 meters.
NYUv2 [60] consists of 24,231 synchronized $640\times 480$ RGB images and depth maps for indoors scenes (household, offices, commercial areas), captured with a Microsoft Kinect. The official split consists of 249 training and 215 test scenes. We use the official test set. Following [3, 88], we remove samples without valid ground truth, leaving 652 valid images for testing. We evaluate on methods on NYU for depth range between $1\times 10^{-3}$ to 10 meters.
SUN-RGBD [62] is an indoor dataset consisting of around 10K images with high scene diversity collected with four different sensors. We use this dataset only for zero-shot evaluation of baselines and our model on the official test set of 5050 images. Evaluation is done on depth values up to 10 meters. Note that we do not use SUN-RGBD for training.
DDAD [29] comprise of diverse dataset of urban, highway, and residential scenes curated from a global fleet of self-driving cars. It contains 17,050 training and 4,150 evaluation frames with ground-truth depth maps generated from dense LiDAR measurements using the Luminar-H2 sensor. We use this dataset only for zero-shot evaluation of baselines and our model, where evaluations are done on depth values up to 80 meters. Note that we do not use DDAD for training.
C Prompts for Natural Text Generation
To generate natural, free-form text that does not follow fixed templates and more closely resembles human descriptions, we use two visual question answering models LLaVA v1.6 Vicuna and LLaVA v1.6 Mistral [47]. To produce diverse, natural captions, we use five different prompts per model. These five prompts are listed in Table 9.
Prompts "Describe the image in one sentence." "Provide a one-sentence description of the image, pay to attention object type." "Capture the essence of the image in a single sentence, pay attention to object relationship." "Condense the image description into one sentence, pay attention to object size." "Express the image in just one sentence, pay attention to the overall layout."
Table 9: Prompts for natural text generation.We use LLaVA v1.6 Vicuna and LLaVA v1.6 Mistral with those 5 prompts to generate 10 sentences of different natural text that adhere to human input for each image.
D Discussion
A single image does not afford metric depth estimation [3, 25, 26, 37, 65, 70, 75, 103]. While some methods [84, 57, 34, 56] opt to predict scaleless relative depth, our method offers a flexible alternative to using additional sensors, e.g., lidar [15, 48, 53, 71, 72, 73, 76, 79], radar [20, 45, 49, 50, 61], or multiple cameras, e.g., stereo [2, 8, 11, 66, 74, 80]. This is facilitated by the the presence of certain objects that tend to co-occur with certain scenes, which can be captured within text descriptions. As 3D scenes are continuous, a short description of objects populating the scene is sufficient to transfer relative depth estimates to metric scale.
Limitations. Our method makes the assumption that there exists an unknown scale in the estimated depth, as modeled by the parameters of a linear transformation. However, this is not always the case as the estimated depth may contain errors, and one may benefit from refinement of the relative depth maps, e.g., [4, 89]. To exploit the invariants of language in the context of depth estimation, future directions may include region-based or even pixel-wise transformations through the use of text descriptions. Also, our method has the advantage of allowing flexible descriptions as input for grounding depth estimate to metric scale. While it offers controllability of 3D reconstructions, it also opens our method to mis-use; malicious users may choose to provide adversarial descriptions to steer predictions incorrectly. Additionally, barring malicious behaviors, uninformative captions regarding the 3D scene at hand, may also yield erroneous transformations.
E RSA Model Architecture
Detailed architecture of RSA model is shown in Table 10. Given text descriptions $\textbf{t}=\{t_{1},t_{2},...\}$, we first encode them into text embeddings and feed them into a 5-layer shared multi-layer perceptron (MLP) to project them into $k=256$ hidden dimensions followed by two separate sets of 5-layer MLPs, one serves as the output head $\psi_{\hat{\alpha}}:\mathbb{R}^{k}\mapsto\mathbb{R}_{+}$ for scale parameter $\hat{\alpha}$ and the other as the output head $\psi_{\hat{\beta}}:\mathbb{R}^{k}\mapsto\mathbb{R}_{+}$ for shift $\hat{\beta}$ parameter. Scale and shift are passed through an exponential function so that they are assumed to be positive in favor of optimization.
Sub-network Layer Units Activation
scene_feat_net Linear 1024 $\rightarrow$ 512 -
LeakyReLU - LeakyReLU
Linear 512 $\rightarrow$ 512 -
LeakyReLU - LeakyReLU
Linear 512 $\rightarrow$ 512 -
LeakyReLU - LeakyReLU
Linear 512 $\rightarrow$ 256 -
LeakyReLU - LeakyReLU
Linear 256 $\rightarrow$ 256 -
shift_net Linear 256 $\rightarrow$ 256 -
LeakyReLU - LeakyReLU
Linear 256 $\rightarrow$ 128 -
LeakyReLU - LeakyReLU
Linear 128 $\rightarrow$ 128 -
LeakyReLU - LeakyReLU
Linear 128 $\rightarrow$ 64 -
LeakyReLU - LeakyReLU
Linear 64 $\rightarrow$ 1 -
scale_net Linear 256 $\rightarrow$ 256 -
LeakyReLU - LeakyReLU
Linear 256 $\rightarrow$ 128 -
LeakyReLU - LeakyReLU
Linear 128 $\rightarrow$ 128 -
LeakyReLU - LeakyReLU
Linear 128 $\rightarrow$ 64 -
LeakyReLU - LeakyReLU
Linear 64 $\rightarrow$ 1 -
Table 10: Structure of RSA Model. Multi-Layer Perceptron: Layers, Units, and Activation Functions.